The length of a rectangle is 3 times the width. If the length is decreased by 4 and width is increased by 1m, perimeter will be 66 m. Find the dimensions of the original rectangle.
Answer by mananth(12270)

width be x
length is 3x
new dimensions
width x+1
length 3x-4
Perimeter = 2*(L+W)
2*(x+1+3x-4)=66
2*(4x-3)=66
8x-6 = 66
add 6 to both sides
8x-6+6= 66+6
8x=72
divide by 8
x= 9 the width
length = 3x = 27